A company's costs, in millions of dollars, are given by the equation. C=x^2-3x-27, where x is the number of items sold, in thousands. What are the costs when 1,000 items are sold? 1,500 items?
I tried substitution 1,000 and 1,500 into the given equation but I was told that was wrong. I do not know where to begin with this problem. HELP PLEASE!!

Answer by checkley79(3341)

X IS REPRESENTED IN 1,000.
THUS YOU ENTER THE AMOUNT IN 1,000.
X=1,000
C=x^2-3x-27
C=1^2-3*1-27
C=1-3-27
C=-29 ANS.
X=1,500
C=1.5^2-3*1.5-27
C=2.25-3-27
C=-27.75 ANS.
######################
ABOUT 7,000 ITEMS MUST BE SOLD TO MAKE MONEY.
C=7^2-3*7-27
C=49-21-27
C=49-48
C=1